At a recent seminar you attended, the invited speaker was discussing some of the advantages and disadvantages of standard costs in terms of evaluating performance and motivating goal-congruent behavior on the part of employees. One criticism of standard costs in particular caught your attention: The use of conventional standard costs may not provide appropriate incentives for improvements needed to compete effectively with world-class organizations. The speaker then discussed so-called continuous-improvement standard costs. Such standards embody systematically lower costs over time. For example, on a monthly basis, it might be appropriate to budget a 1.0% reduction in per-unit direct labor cost. Assume that the standard wage rate into the foreseeable future is $30.00 per hour. Assume, too, that the budgeted labor-hour standard for October of the current year is 1.50 hours and that this standard is reduced each month by

1.0%. During December of the current year the company produced 10,000 units of XL-10, using 14,800 direct labor hours. The actual wage rate per hour in December was $32.00. Required: 1. Prepare a table that contains the standard labor-hour requirement per unit and standard direct labor cost per unit for the 4 months, October through January. 2. Compute the direct labor efficiency variance for December. Was this variance favorable (F) or unfavorable (U)?

  • Required 1
  • Required 2

Prepare a table that contains the standard labor-hour requirement per unit and standard direct labor cost per unit for the 4 months, October through January. (Do not round intermediate calculations. Use rounded answers in the subsequent requirement. Round your "Standard Direct Labor Cost/Unit" answers to 2 decimal places and "Standard Labor-Hour Requirement/Unit" answers to 5 decimal places.)

Month Standard Labor-Hour Requirement/Unit Standard Direct Labor Cost/Unit
October hr.per unit
November hr.per unit
December hr.per unit
January hr.per unit

Compute the direct labor efficiency variance for December. Was this variance favorable (F) or unfavorable (U)? (Do not round intermediate calculations. Round your final answer to nearest whole dollar amount.)

Direct-labor efficiency variance
Answer #1

1.

Month standard labour hour requirement / unit standard direct labour cost / unit
October 1.50 $30
November 1.50-1% = 1.485 30-1% = 29.7
December 1.485-1% = 1.47015 29.7-1%=29.40
January 1.47015-1%=1.45544 29.40-1%=29.10

2.december month

Standard hours required for actual output = 10,000 × 1.47015

(Sh)=14,701.5

Standard rate (sr) = 29.40

Actual hours (ah) = 14,800

Labour efficiency variance = (sh - ah ) × sr

= ( 14,701.5 - 14800 ) × 29.40

= ($ 2896) adverse